OpenStudy (anonymous):

My Flower Basket delivers flowers within a 20 mile radius of their store. If this area is represented graphically, with My Flower Basket located at (5, 2), what is the equation that represents the delivery area?

OpenStudy (anonymous):

A (x - 2)2 + (y - 5)2 = 20 B (x - 5)2 + (y - 2)2 = 20 C (x + 5)2 + (y + 2)2 = 400 D (x - 5)2 + (y - 2)2 = 400

OpenStudy (anonymous):

(x – 5)2 + (y – 2)2 = 400

OpenStudy (anonymous):

d is your answer !

OpenStudy (anonymous):

Thanks hunn (: you sure?

OpenStudy (anonymous):

welcome and yes im sure lol (x – h)2 + (y – k)2 = r2 <---- is the equation of a circle and 20^2=400

OpenStudy (anonymous):

Thankk you!! <3

OpenStudy (anonymous):

welcome XD